# Evaluating Open Language Models Across Task Types, Application Domains, and Reasoning Types: An In-Depth Experimental Analysis

### v1

@@ 3.3 Comparison Across Application Domains
Particularly for pre-trained models, the performance is very sensitive across domains. For social sciences & humanities, and science & technology domain groups, Falcon-2-11B performs the best with Gemma-2B and Llama-3-8B following. Gemma-2B and Falcon-2-11B are not always best. In health and medical tasks, Gemma-7B performs the best. Therefore, for domains, the choice of pre-trained LMs significantly depends on the use case and other constraints.
Among the IT models, we see similar trends as in task types - Mistral-7B-I being the best in all domains, and Gemma-2B-I coming second for all domains. The difference is closer in some domains like Computer Science, News, and Books, and largest in Economics. We don’t observe any patterns in differences along domain groups in this case. Phi-3-mini-128k is still sensitive to domains and under-performing. Hence, Mistral-7B-I is still the best choice with best prompt style if the available resource allows, and if not, then Gemma-2B-I is the way to proceed.
@@ 3.4 Comparison Across Reasoning Types
In the pre-trained models, we find that where reasoning is involved, Gemma-2B marginally outperforms Falcon-2-11B in all types of reasoning except Abductive reasoning, where it comes second by a small margin. Llama-3-8B proves to be the best in analogical reasoning. Therefore, where reasoning is involved, Gemma-2B becomes a preferred choice over Falcon-2-11B, having advantage in both performance and model size.
With the IT models, the behavior here remains similar to the previous two aspects for all the five models. The difference between Mistral-7B-I and Gemma-2B-I is minimum in complex inference & analysis types, and maximum for types like logical and quantitative reasoning. This shows that while choosing a pre-trained model has its complexities, for IT models, the choice is relatively simpler after considering external constraints.
The detailed quantified performance of each entity of all three aspects (even those not included in Figure 3, but available in the dataset) with each LM is given in Appendix B.1.
@@ 3.5 Comparison Across Prompt Styles
We visualize the results using a line graph shown in Figure 6. Using this, users can analyze the trade-offs of crafting instructions versus its possible impact on performance.
On initial analysis, using definitions seems better for all the tasks, and the performance increase looks small after 2 examples. So, even if 8 examples is optimal, using two might suffice. This trend is consistent for most entities across all three aspects.
However, adding definition impacts different entities differently. For example, dialogue act recognition’s performance on zero examples increases from 80.37 to 88.77 just by including task description. But, for keyword tagging, the change is from 82.73 to only 82.81. We also see behaviors like Word Analogy, for which in absence of definition, more examples negatively impact the output. It may be because in absence of clear instruction, the model fails to comprehend the task from examples.
Among the application domains, taking ‘Social Media’, adding task definition increases performance from 82.27 to 91.58 without examples, but, if difficult, adding 2 examples without definition also improves score to 93.17. So, a choice is available between definition and examples. The rate of improvement with adding examples is also different for different entities. Note that, some tasks don’t have 8 examples in the dataset, so 4 to 8 example transition should be inferred accordingly.
It is not possible to discuss each of the trends, but using these graphs, one can determine a good instruction template for an application within other constraints of ability, cost, need, etc. constraints of crafting instruction.
These trends are different for different LMs. So, we have included these line graphs for all other LMs in Appendix E.2. This will also help in analyzing best prompt style and studying relative performance difference of each entity of each aspect.
@@ 3.6 Task Definition v/s Paraphrased Definition
To evaluate how sensitive the models are to the provided task definition, we also evaluate them using paraphrases of task definitions. These are generated using gpt-3.5-turbo (Brown et al., 2020; OpenAI, 2023), and used with in-context examples as per Table 1. The results are generated and evaluated using the same pipeline. The outcome of this experiment is given in Table 2 for the two-best performing LMs in each category.
@@
The median decrease in performance for all 10 LMs also is only 0.35%., which can be attributed to some loss of information during paraphrasing. But, most of the models prove robust to perturbations in task definitions, as long as a prompt can reasonably explain the task. One exception was Phi-3-mini-128k-I, for which the performance degraded by 33.64%, making it very sensitive to prompt. Appendix E.3 has more details on obtaining paraphrases and results on all LMs.
3.7 Comparison with State-of-the-art
(a) Task Types
(b) Domains
(c) Reasonings
We compare models used in our experiments with GPT-3.5-Turbo (GPT-3.5-T), GPT-4o (OpenAI et al., 2024; OpenAI, 2023) – OpenAI’s most recent model, and DeepSeek-v2 (DS-2) (DeepSeek-AI, 2024). These models give SOTA performance, but have constraints discussed previously. GPT-3.5-Turbo is a 175B model, and GPT-4o is reported to be larger. Both are closed models accessible only by APIs. GPT-4o is also very costly. DeepSeek-v2 is open, but has 243B parameters. We compare Gemma-2B, Falcon-11B, Mistral-7B-I and Gemma-2B-I here, as they to give best results. We use 8 examples with task definition prompt for SOTA models, and report variation across aspects in Figure 7.
LM DS-2 GPT-3.5-T GPT-4o
Gemma-2B 4.39% 8.63% 9.78%
Falcon-2-11B 5.19% 8.87% 10.02%
Gemma-2B-I 3.05% 6.89% 8.06%
Mistral-7B-I -3.17% 0.86% 2.11%
Table 3: Average Percentage decrease in mean BERTScore recall of various language models compared to DeepSeek-v2 (DS-2), GPT-3.5-Turbo and GPT-4o.
We can see that Mistral-7B-I matches closely with all three models across most entities of all aspects, and even beats DeepSeek-v2 and GPT 3.5-Turbo occasionally. Gemma-2B-I also competes closely with DeepSeek-v2. From the average % decrease in performance reported in Table 3, Mistral-7B-I performs better than DeepSeek-v2, and is within 2% of GPT-4o. Gemma-2B, Gemma-2B-I also are relatively close considering they are approx 100-fold smaller. The gaps of pre-trained models are higher than IT models, but we believe aligning them further for specific use can improve responses. We also evaluate these SOTA LMs for all entities of each aspect in Appendix B.1.
4 Key Takeaways
@@
We find that recent, open and small-scale Language Models (LMs) are very effective. No one LM is a global solution, but if selected and used appropriately for a task type, domain, or reasoning type, they can perform within acceptable levels large SOTA LMs while giving other advantages of efficiency, control and cost-effectiveness.
@@
Among pre-trained models, this variation is more visible, with Gemma-2B and Falcon-2-11B performing best on many aspects and entities, but sometimes, they are weak and Gemma-7B, Llama-3-8B perform better, particularly for some application domains (also detailed in Appendix B.1). The performance of pre-trained models can be taken as a measure of their knowledge of different entities. Based on other factors like availability, compliance, size, appropriate LMs can be selected and customized using fine-tuning, prompting or in a custom pipeline for further use. Limitations of some pre-trained models are discussed in Appendix C.2.
@@
For IT models, we see that Mistral-7B-I is the strongest in all aspects, and Gemma-2B-I comes second best in most cases. Since these models are IT, they can also be used directly with chat-style description and examples. Between these two (and other models), based on other factors, the right choice can be made. Some qualitative outputs of Mistral-7B-I are given in Appendix C.1.
@@
Our experimentation also includes analyzing the performance trade-off for multiple prompt styles and recommending the best one for these models. As noted in Section 3.5 and Table 1, having a chat-style task description to guide the LM is generally more helpful. But, having more in-context examples is not always better, and considering use-case and LMs, the right number of example can vary. The models are also robust to changes in task definitions, if it can provide all (or most) information to complete the task. They are also reasonably robust to subtle intended/unintended incorrectness in definitions, which is analyzed in Appendix E.4.
When comparing with SOTA, these models are capable of matching (and sometimes even outperforming) large models like DeepSeek-v2, GPT-3.5-Turbo and GPT-4o within 10% decrease in performance. This shows that if the right LM is selected, one can utilize these models without significant performance degradation while benefiting from aforementioned advantages.
@@ 5 Conclusion
In this study, we identify some limitations of using large-scale language models and show that open LMs with 2B–11B parameters can be effective for applications if selected properly. We use a dataset derived from Super-Natural Instructions (Wang et al., 2022) and analyze semantic correctness of output on 12 task types, 12 application domains and 10 reasoning types. Despite being 10-20 folds smaller, these models are effective and can compete with (and sometimes outperform) models like GPT-3.5-Turbo, DeepSeek-v2 and GPT-4o. We provide a guide in Appendix A on how to use this work to select an LM for one’s specific needs. We hope that our analysis will enable the community to make an informed selection of LM and prompt that best meet their needs.
@@ 6 Limitations
Limitations of Dataset: We derive our experimental dataset from the test-set of Super Natural Instructions (Wang et al., 2022) and take the labels of aspects from there itself. We therefore assume that they are labeled correctly for task types, domains and reasoning types. There are many other task types, domains and reasoning types not available in its test set, which we were not able to consider. However, if an entity of an aspect is not present, one can leverage the performance of the groups that we created in Figure 3, or choose the nearest entity (from Section 3/ Appendix B.1) and roughly estimate the performance. We were also bounded in prompts by the examples and task definitions available. But, we did experiment by paraphrasing definitions in Section 3.6, Appendix E.3 to limit this to some extent.
@@
Limitations of LMs: There are many LMs proposed by the research community, and it is not practically feasible to experiment with all of them. Further, the definition of a small LM is also relative. We selected the LMs based on the ones which have come out recently and promise strong capabilities. Although we capped our LMs at 11B parameters, we didn’t find the performance to be a function of number of parameters, so we believe this decision should not have too drastic effects. We also didn’t evaluate biases (Gallegos et al., 2024) and other factors other than semantic correctness of generated results of these models, but the models’ technical reports (Abdin et al., 2024; Jiang et al., 2023; Team et al., 2024; AI@Meta, 2024; Almazrouei et al., 2023; TIIUAE, 2024) provide more details on those.
@@
Limitations of Prompts: We experimented with 8 prompt styles, apart from using paraphrased definitions, adversarial definitions. But, all of them were built using the elements of the dataset available. We acknowledge that there may be some tasks where another prompt style or using more domain-adapted prompts perform better. Additionally, if the LM is adapted/fine-tuned in any way, the best prompt style can change based on the data and technique used for it. However, this study should provide an initial idea of whether descriptions are needed and the number of examples required when using the LM without any changes.
@@
Assumptions in Reporting of Results: We are considering the impact of each aspect one at a time when reporting results. For example, in pre-trained models, we see that Gemma-2B is best for Grammar Error Correction, and Falcon-2-11B is best for Economics domain. But what if there is a task instance that involves grammar error correction for an Economics article? This can sometimes give a dual outcome (not always). To eliminate this, we tried to do a pairwise aspect analysis, but in the dataset, 86.86% of task type-domain and 88.25% of domain-reasoning type pairs had no task instances. We could have generated labels of aspect entities using other techniques, or could’ve generated artificial data to fill these gaps, but we didn’t want conflicting sources of experimental data as they could created additional undesired biases/variations of sources, type of data. Therefore, due to lack of sufficient labels, we didn’t report those results. However, despite this independent assumption, this work can still help in narrowing down to 2-3 models which will be helpful. We also assume that the results reported by our experimental data represents the actual performance of that LM on that entity type. This may not be correct but considering Super Natural Instructions to be a meta-dataset of many other datasets, we believe it is a reasonable assumption.
@@ Appendix A Guide to LM Selection for Your Application Using this Work
Then, check the relative performance of LMs for your task type/domain/reasoning type (or a combination). Find the closest available entity, and look up the performance of LMs of interest from Tables 576. From there, consider some options based on other constraints. For example, if you are planning to fine-tune LMs on your task, choose from pre-trained models, if not, utilizing IT models will be better, or if you are bounded by resources, consider using smaller models that fit the requirements.
@@
This work is accompanied by an anonymous GitHub repository  [^3] https://github.com/neelabhsinha/lm-application-eval-kit as a utility which will allow evaluating any LM as per this framework and generating visualizations. It supports evaluation and generation of visualizations on other evaluation metrics that are discussed in Table 1, and on a different set of task types, application domain and reasoning types as needed with minor configuration changes. No code change will be needed for utilizing HuggingFace implemented models. Usage guidelines are available in the README of the repository.

### v2

@@ 3.3 Comparison Across Application Domains
Particularly for pre-trained models, the performance is very sensitive across domains. For social sciences & humanities, and science & technology domain groups, Falcon-2-11B performs the best with Gemma-2B and Llama-3-8B following. Gemma-2B and Falcon-2-11B are not always the best ones. In health and medical tasks, Gemma-7B outperforms all models. Falcon-2-11B and Gemma-2B suffer a significant performance degradation in this group. Therefore, for domains, the choice of pre-trained LMs depends on the use case and other constraints. SmolLM-1.7B felt like a strong choice in task types, but here we see here that it struggles with these domains. It’s strength in Section 3.2 might be from other domains not considered here, showing its sensitivity with domains.
Among the IT models, we see similar trends as in task types - Mistral-7B-I being the best in all domains, and Gemma-2B-I and SmolLM-1.7B-I competing for second. The difference with Gemma-2B-I is closer in some domains like Computer Science, News, and Books, and largest in Economics. We also see that SmolLM-1.7B-I has strong limitations in Science and Technology group. Hence, Mistral-7B-I is still the best choice with best prompt style if the available resource allows, and if not, then Gemma-2B-I or SmolLM-1.7B is the way to proceed based on requirements.
Group-level behavior is more prominent in this aspect, highlighting the importance of our three-tier framework. Even in case of analyzing a new domain that is not present here, the performance of the group that domain would belong to can give an idea of baseline performance.
@@ 3.4 Comparison Across Reasoning Types
In the pre-trained models, we find that where reasoning is involved, Gemma-2B marginally outperforms Falcon-2-11B in all types of reasoning except Abductive reasoning, where it comes second by a small margin. It shows that Gemma-2B is a great choice where reasoning is involved, having advantage in both performance and model size. Llama-3-8B proves to be the best in analogical reasoning. In general, it is observed that the performance of all pre-trained LMs is the least for Comparative and Relational reasoning types, highlighting a potential common limitation of ability in these types of task in zero-shot. Therefore, adapting the LMs might become crucial in this case.
With IT models, behavior remains similar to the previous two aspects for all the five models, with Mistral-7B-I coming out to be a clear choice. The difference between Mistral-7B-I and Gemma-2B-I is minimum in complex inference & analysis types, and maximum for types like logical and quantitative reasoning. SmolLM-1.7B-I also depicts weaknesses in some reasoning types. This shows that while choosing a pre-trained model has its complexities, for IT models, the choice is relatively simpler after considering external constraints.
The quantified performance of each entity of all three aspects in the dataset (even ones not included in Fig 3) with each LM is given in Appendix B.
3.5 Comparison with State-of-the-art LLMs
(a) Task Types
(b) Domains
(c) Reasonings
We compare small, open LMs Gemma-2B, Falcon-11B, Mistral-7B-I and Gemma-2B-I (overall best two from each category) with recent SOTA LLMs like GPT-4o-mini, GPT-4o (OpenAI et al., 2024; OpenAI, 2023), DeepSeek-v2 (DS-2) (DeepSeek-AI, 2024) and Gemini-1.5-Pro (Team et al., 2024a). GPT-4o, Gemini-1.5-Pro and GPT-4o-mini are costly, large, closed models accessible using APIs. DeepSeek-v2 is open, but has 243B parameters. We use 8 examples with task definition for SOTA models, and report results in Figure 6.
We witness that Mistral-7B-I matches closely with all SOTA models globally. It’s even very close to GPT-4o in some groups like Generation tasks, Art and Literature, and Media and Entertainment domains. All the 4 models outperform GPT-4o-mini, Gemini-1.5-Pro and DS-2 in many categories where they are strong, proving them to be a very strong choice. In application domains like in Social Sciences and Humanities group and Art and Literature group, Gemma-2B and Gemma-2B-I outperform Gemini-1.5-Pro as well. Being the open-sourced variant of a close family, this is commendable and shows that open LMs can be better choices than large or expensive ones in some usage scenarios. Many inferences can be drawn from the graph based a reader’s need through this evaluation framework. From the average global % decrease in performance reported in Table 1, these models are globally competitive with the SOTA LLMs, proving their readiness in being utilized for practical applications with their other advantages as discussed previously. The gaps of pre-trained models are higher than IT models, but aligning them further for specific use can improve results. We also evaluate these SOTA LMs for all entities of each aspect in Appendix B.
LM Gem-1.5 G-4o-m G-4o
Gemma-2B 3.28% 8.12% 9.78%
Falcon-2-11B 3.54% 8.37% 10.02%
Gemma-2B-I 1.44% 6.38% 8.07%
Mistral-7B-I -4.94% 0.32% 2.12%
Table 1: Avg. Percentage decrease in mean BERTScore recall of open LMs compared to Gemini-1.5-Pro (Gem-1.5), GPT-4o-mini (G-4o-m) and GPT-4o (G-4o).
@@ 3.6 Comparison Across Prompt Styles
We visualize the results in Figure 7. Using this, users can analyze the trade-offs of crafting instructions versus its possible impact on performance.
On initial analysis, using chat-style definitions proves better, but the performance increase looks small after 2 examples. So, using 2 examples can suffice. This trend is consistent for most entities across all three aspects. However, adding definition impacts different entities differently. For example, dialogue act recognition’s performance on zero examples increases from 80.37 to 88.77 just by including task definition. But, for keyword tagging, the change is from 82.73 to only 82.81. We also see behaviors like Word Analogy, for which more examples negatively impact the output if definition is not provided. It may be because in absence of clear instruction, the model fails to comprehend the task from examples. Further, taking ‘Social Media’, adding task definition increases performance from 82.27 to 91.58 without examples, but, adding 2 examples without definition also improves score to 93.17. So, a choice is available between definition and examples. The rate of improvement with adding examples is also different for different entities. Some tasks don’t have 8 examples in the dataset, so 4 to 8 example transition should be inferred accordingly.
Using these graphs, one can determine a prompt style for an application within other constraints of ability, cost, need, etc. in crafting instructions. These trends are different for different LMs. So, we have included these line graphs for all other LMs in Appendix D.2. This will also help in analyzing best prompt style and studying relative performance difference of each entity of each aspect.
@@ 3.7 Task Definition v/s Paraphrased Definition
To evaluate dependency of models to the provided task definition, we also evaluate them with their paraphrases. These are generated using gpt-3.5-turbo (Brown et al., 2020; OpenAI, 2023), and used with best in-context example count as per Table 7. Then, results are evaluated using the same pipeline, and reported in Table 2 for the two-best performing LMs in each category.
@@
The median decrease in performance for all 10 LMs also is only 0.35%., which can be attributed to some loss of information during paraphrasing. But, most of the models prove robust to perturbations in task definitions, as long as a prompt can reasonably explain the task. Appendix D.3 has more details on obtaining paraphrases and results on all LMs.
4 High-level Takeaways
@@
We find that recent, open and small-scale Language Models (LMs) are very effective. Detailed recommendations on LMs and their performance trends in different groups and entities are discussed in depth in Sections 3.23.3 and 3.4, but we summarize them in the below paragraphs too. Although it is visible that no single LM is a global solution, but, if selected and used appropriately with an effective prompt style for a task type, domain, or reasoning type, they can perform within 10% (worst case) of SOTA LLMs like GPT-4o, and outperform DS-2, GPT-4o-mini, and Gemini-1.5-Pro with advantages in efficiency, control and cost.
@@
For the LMs we experimented with, among pre-trained models, we recommend using Gemma-2B and Falcon-2-11B based on different aspects and entities, but sometimes, Gemma-7B, Llama-3-8B can be great for some application domains (also detailed in Appendix B). The performance of pre-trained models can be taken as a measure of their knowledge of different use-cases. Based on other factors like availability, compliance, size, right LM can be selected and customized as needed. Limitations of some pre-trained models are discussed in Appendix F.2.
@@
For IT models, Mistral-7B-I is a clear best in all aspects, and Gemma-2B-I and SmolLM-1.7B-I come second in most cases. Since these models are IT, they can be used directly with chat-style description and examples. We recommend a model in these three (and other models), based on other factors like size, licensing, etc. Some qualitative outputs of Mistral-7B-I are given in Appendix F.1.
@@
We also study the performance trade-off for multiple prompt styles and recommend the best one for these models. As noted in Section 3.6, having a chat-style task description to guide the LM is recommended. But, having more in-context examples is not always better, and considering use-case and LMs, the right number of example can vary. The models are also robust to changes in task definitions, if it can provide all (or most) information to complete the task. They are also reasonably robust to subtle intended/unintended incorrectness in definitions, which is analyzed in Appendix D.4.
@@ 5 Conclusion
We identify some limitations of using SOTA, proprietary LLMs and show that open LMs with 1.7B–11B parameters can be effective for applications. We create a three-tier evaluation framework and analyze semantic correctness of output of 10 LMs across multiple hierarchical umbrellas. Using this framework, we demonstrate that while these models don’t work best in every scenario, if selected properly, they are effective and can compete with and outperform models like Gemini-1.5-Pro, DeepSeek-v2 and GPT-4o-mini and GPT-4o despite being 10-20 folds smaller in size. We also provide a guide in Appendix A on how one can this work to select an LM for one’s specific needs. We hope that our contributions will enable the community to make a confident shift towards considering using these small, open LMs for their need.
@@ 6 Limitations
Limitations of Dataset: We derive our experimental dataset from the test-set of Super Natural Instructions (Wang et al., 2022) and take the labels of aspects from there itself. We therefore assume that they are labeled correctly for task types, domains and reasoning types. There are many other task types, domains and reasoning types not available in its test set, which we were not able to consider. However, if an entity of an aspect is not present, one can leverage the performance of the groups that we created in Figure 3, or choose the nearest entity (from Section 3/ Appendix B) and roughly estimate the performance. We were also bounded in prompts by the examples and task definitions available. But, we did experiment by paraphrasing definitions in Section 3.7, Appendix D.3 to limit this to some extent. Using this dataset however may not bring significant dataset bias, as this is a meta-dataset curated using several NLP benchmark datasets.
@@
Limitations of LMs: There are many LMs proposed by the research community, and it is not practically feasible to experiment with all of them. Further, the definition of a small LM is also relative. We selected the LMs based on the ones which have come out recently and promise strong capabilities. Although we capped our LMs at 11B parameters, we didn’t find the performance to be a direct function of number of parameters, so we believe this decision should not have too drastic effects. We also didn’t evaluate biases (Gallegos et al., 2024) and other factors other than semantic correctness of generated results of these models, but the models’ technical reports (Allal et al., 2024; Team et al., 2024b; AI@Meta, 2024; Almazrouei et al., 2023; TIIUAE, 2024) provide more details on those.
@@
Limitations of Prompts: We experimented with 8 prompt styles, apart from using paraphrased definitions, adversarial definitions. But, all of them (excluding paraphrased and adversarial definitions) were built using the elements of the dataset available. We acknowledge that there may be some tasks where another prompt style or using more domain-adapted prompts perform better. Additionally, if the LM is adapted/fine-tuned in any way, the best prompt style can change based on the data and technique used for it. However, to keep a standard and common features across tasks, we intentionally chose this approach. This study should provide an initial idea of whether descriptions are needed and the number of examples required when using the LM without any changes.
@@
Assumptions in Reporting of Results: We are considering the impact of each aspect one at a time when reporting results. For example, in pre-trained models, we see that Gemma-2B is best for Grammar Error Correction, and Falcon-2-11B is best for Economics domain. But what if there is a task instance that involves grammar error correction for an Economics article? This can sometimes give a dual outcome, with one LM recommended for task type (grammar error correction), and one for domain (Economics). To eliminate this, we tried to do a pairwise aspect analysis, but in the dataset, 86.86% of task type-domain and 88.25% of domain-reasoning type pairs had no task instances. We could have generated labels of aspect entities using other techniques, or could’ve generated artificial data to fill these gaps, but we didn’t want conflicting sources of experimental data as they could created additional undesired biases/variations of sources, type of data. Therefore, due to lack of sufficient labels, we didn’t report those results. However, despite this independent assumption, this work can still help in narrowing down to 2-3 models which will be helpful. We also assume that the results reported by our experimental data represents the actual performance of that LM on that entity type. This may not be correct but considering Super Natural Instructions to be a meta-dataset of many other datasets, we believe it is a reasonable assumption.
@@ Appendix A Guide to LM Selection for Your Application Using this Work
Then, check the relative performance of LMs for your task type/domain/reasoning type (or a combination). Find the closest available entity, and look up the performance of LMs of interest from Tables 465. From there, consider some options based on other constraints. For example, if you are planning to further align LMs on your task using any technique, choose from pre-trained models, if not, utilizing IT models will likely yeild better results. If you are bounded by resources, consider using smaller models that fit the requirements, or if you are bound by business/regulatory constraints, choose accordingly.
@@
This work is accompanied by a GitHub repository linked in the first page of the paper as a utility which will allow evaluating any LM as per this framework and generating visualizations. It supports evaluation and generation of visualizations on other evaluation metrics that are discussed in Table 7, and on a different set of task types, application domain and reasoning types as needed with minor configuration changes. No code change will be needed for utilizing HuggingFace implemented models. Usage guidelines are available in the README of the repository.